# Interacting-cluster spin liquids with robust flat bands evolving into higher-rank half-moon phases and topological Lifshitz transitions

### v2

@@ Abstract
Classical spin liquids are disordered magnetic phases, governed by local constraints that often give rise to flat-band ground states. When constraints take the form of a zero-divergence field within a cluster of spins, the spin liquid is often described by an emergent Coulomb gauge theory. Here we introduce an interaction $\eta$ between these clusters of spins which compete with the zero-divergence field. Using a framework embracing both the connectivity matrices of graph theory and the topology of band structures, we develop a generic theory of interacting-cluster Hamiltonians. We show how flat bands remain at zero energy up to finite interaction $\eta$, until a dispersive band becomes negative, stabilizing a spiral spin liquid with a hypersurface of ground-state manifold in reciprocal space. This hypersurface serves as a mold for the apparition of the half-moon patterns in the equal-time structure factor. Our generic approach enables to extend the notion of half moons to the perturbation of higher-rank Coulomb fields and pinch-line spin liquids. In particular, multi-fold half moons appear when unconventional gauge charges, such as potential fractons, are stabilized in the ground state. Finally, half-moon phases can be tuned across the equivalent of a Lifshitz transition, when the hypersurface manifold changes topology.
@@ I Introduction
Using tools recently developed for the classification of classical spin liquids [20, 21, 22, 23, 24] together with graph-theory methods based on the connectivity matrix [25, 26, 10, 13], we build a generic theory of interacting-cluster systems for a seemingly infinite variety of lattices and clusters in two and three dimensions (or higher). This interaction $\eta$ creates an energetic competition with the zero-divergence field required by the Coulomb gauge theory, and will be our flat-band engineering parameter. We quantify the stability of the flat-band spin liquid until $\eta$ reaches a critical value, where the energetic competition ultimately brings the lowest dispersive band below the flat bands. This is where pinch points transform into half moons, the signature of a spiral spin liquid where the magnetic texture of gauge charges is stabilized in the ground state. Then we extend the notion of half moons to higher-rank U(1) gauge fields, including models with pinch lines. Since multifold half moons come from the dispersive band of a parent rankn-n Coulomb field, they represent the signature of rankn-n gauge charges, such as fractons, in the ground state. Finally, we show how half-moon phases can be adiabatically tuned across the equivalent of a Lifshitz transition separating distinct spiral spin liquids.
@@ II Cluster Hamiltonians
2D Lattice Kagome Checkerboard Hexagonal
Kagome
hexagonal
Square
octagon
Ruby Square kagome
Decorated
square kagome
Octagonal
kagome
Lattice
Scheme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Premedial
Lattice
Scheme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
3D Lattice Pyrochlore Octahedral
Kagome
bipyramidal
Quadrupahedral Hyperkagome
Lattice
Scheme
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ II.2 Flat-band ground state
From now on, we shall work within the Luttinger-Tisza approximation (LTA) [37, 38, 39], where the spin length constraint is only enforced on average over the entire system, $\sum_{i}|\mathbf{S}_{i}|=N$ with $N$ the total number of spins. A known limitation of the LTA is that it cannot always properly account for thermal order by disorder [40, 27], i.e. when thermal fluctuations entropically lift the classical ground-state degeneracy and favor magnetic order at very low temperatures. It is, however, possible to estimate when order by disorder is expected or not, i.e. when flat bands are expected to lead to a magnetically disordered spin-liquid ground state (see section II.5). And even if the system ultimately orders at very low temperature $T_{c}$, the results of this paper remain valid at temperatures above $T_{c}$. Indeed, even if the LTA is an approximation, this method is particularly well suited to the study of classical spin liquids whose disordered magnetic texture accommodates well the averaged spin length constraint, as confirmed by many comparison between theory and simulations [41, 42, 10, 20, 22, 23]. And since we shall only consider isotropic exchange coupling here, the spin component $\alpha\in\{x,y,z\}$ is actually irrelevant and will be omitted from now on.
@@ II.3 The connectivity matrix
$\begin{split}h^{v\leftarrow c}_{\mu,X}(\mathbf{q})&=\sum_{j}h^{v\leftarrow c}_{(0,\mu),(j,X)}e^{i(-\mathbf{R}_{j}+\mathbf{r}_{\mu}-\mathbf{r}_{X}^{c})\cdot\mathbf{q}}\\ &=\sum_{k}h^{v\leftarrow c}_{(k,\mu),(0,X)}e^{i(\mathbf{R}_{k}+\mathbf{r}_{\mu}-\mathbf{r}_{X}^{c})\cdot\mathbf{q}}\\ &=\left[h^{c\leftarrow v}(\mathbf{q})\right]^{\dagger}\end{split}$ (12)
@@ IV.1 General theory of half moon patterns on uniform interacting-cluster systems
Since $\lambda_{ld}(\mathbf{q}^{\star})=0$ by definition of the the band-touching point $\mathbf{q}^{\star}$, $\mathbf{q}^{\star}$ remains a band-touching point for all values of $\zeta$, even when the dispersive band $\Lambda_{ld}$ becomes negative. By analytical continuity of Eq. (37), the manifold $\mathcal{Q}$ typically forms a closed contour in 2D or a closed surface in 3D, in other words a hypersurface, around $\mathbf{q}^{\star}$ in reciprocal space.
@@
In absence of extensively degenerate flat bands, this half-moon phase is not a traditional spin liquid. But since the manifold $\mathcal{Q}$ is sub-extensive and isomorphic to an (d1)(d-1)-sphere within a given Brillouin zone, the half-moon phase corresponds to a so-called spiral spin liquid [56, 57, 58, 59, 60, 61] with unusual, non-local, spin dynamics [62]. At very low temperature, beyond the Luttinger–Tisza approximation, thermal order-by-disorder (ObD) might entropically lift the degeneracy of manifold $\mathcal{Q}$, especially in $d=3$ dimensions. For most models in $d=2$ dimensions though, the Hohenberg-Mermin-Wagner theorem should prevent magnetic order, even if the breaking of discrete lattice rotational symmetry remains a possibility [63].
In any case, for $\zeta>\zeta_{c}$, the energy minimum of the interacting-cluster Hamiltonian ($\Lambda_{ld}$) comes from the lowest excited dispersive band of the parent cluster Hamiltonian ($\lambda_{ld}$). Since the flat bands support an emergent Coulomb gauge theory, the lowest dispersive band supports a gapless form of the gauge-charge excitations of the Coulomb theory [10, 9]. The type of flat-band engineering we have introduced via the cluster interaction $\eta$ thus stabilizes magnetic textures where gauge charges have become part of the ground state. At low temperature, fluctuations are restricted to perturbations close to the hypersurface manifold. But at intermediate temperatures, when the flat bands start to be populated, there is a temperature window where a dense phase of gauge charges co-exist with the vacuum of the flat bands. By varying the parameter $\eta$, one can include more or less of the dispersive band below the flat bands, i.e. qualitatively more or less gauge charges in the low-energy regime. For models where the thermal order-by-disorder selects a long-range magnetic order at very low temperatures, the ObD transition would correspond to the crystallization of these gauge charges [64, 10].
@@ IV.5.1 Gapless systems
For models with a single cluster type (e.g. checkerboard, octochlore), one has $\lambda(\mathbf{q})\equiv|\mathbf{L}(\mathbf{q})|^{2}$. If the parent system dispersive band $\lambda(\mathbf{q})$ possesses a saddle point, then, as the threshold $\lambda_{c}(\zeta)$ rises past that saddle energy $E_{s}$ (see Eqs. (52,67)), the topology of the ground–state manifold $\mathcal{Q}$ changes, as illustrated in Fig. 7(b) for the generalized checkerboard lattice. The key point to understand is that the manifold $\mathcal{Q}$ varies continuously with $\zeta$. For small $\zeta$, the manifold $\mathcal{Q}$ encircles the pinch points as expected for half moons; see the blue contours on the weight $W_{2}(\mathbf{q})$ of the dispersive band (right panel of Fig. 7(a)) and the half moons in the structure factor (left panel).
@@ IV.5.2 Gapped systems
$\Leftrightarrow\zeta$ $\geq$ $\zeta_{\Delta}>\zeta_{c}(\Delta)$ (73)
@@ V Discussion
Here we have build a generic theory for multifold half moons, including the evolution of pinch lines into extended surfaces, as the signatures of spiral spin liquids whose magnetic texture contains higher rank gauge charges in the ground state. The manifold of minimum energy forms a closed contour whose topology changes at Lifshitz transitions, when the dispersive band of the parent cluster Hamiltonian possesses saddle points. In most, if not all, three-dimensional models, thermal order by disorder will take place at very low temperature and long-range order the system. As was the case on pyrochlore [10], this magnetic ordering can be seen as the crystallization of the gauge charges that populate the dispersive band. In the case of higher rank spin liquids, one might thus be able to observe a crystallization of fractons. When the parent cluster Hamiltonian is gapped, half moons disappear but spiral spin liquids and Lifshitz transitions remain.
The present framework thus offers an infinite variety of theoretical models supporting a dense phase of gapless gauge charges at intermediate temperature, stabilized by cluster interaction. Their non-local dynamics is an open question, and their connection with local momentum vortices found in spiral spin liquids [62] – equivalent to quadrupoles of fractons – is an appealing direction to investigate. The investigation of half moons in quantum spins $S=1/2$ has given interesting physics on kagome [72, 54, 73] and can now be applied to higher-rank and pinch-line spin liquids.
@@ Appendix B Self Consistent Gaussian Approximation
$\begin{split}\mathcal{H}&=\frac{1}{2}\sum_{i,j}H_{i,j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}\\ &=\frac{1}{2N}\sum_{\mathbf{q}}\sum_{\mu,\nu}H_{\mu\nu}(\mathbf{q})\mathbf{S}_{\mu}(\mathbf{q})\cdot\mathbf{S}_{\nu}(\mathbf{q})\end{split}$ (88)

### v3

@@ Abstract
Classical spin liquids are disordered magnetic phases, governed by local constraints that often give rise to flat-band ground states. When constraints take the form of a zero-divergence field within a cluster of spins, the spin liquid is often described by an emergent Coulomb gauge theory. Here we introduce an interaction $\eta$ between these clusters of spins which compete with the zero-divergence field. Using a framework embracing both the connectivity matrices of graph theory and the topology of band structures, we develop a generic theory of interacting-cluster Hamiltonians. We show how flat bands remain at zero energy up to finite interaction $\eta$, until a dispersive band becomes negative, stabilizing a spiral spin liquid with a hypersurface of ground-state manifold in reciprocal space. This hypersurface can be interpreted as an effective Fermi surface in the spectrum of the parent system, acting as a tunable energy selector despite the absence of particle filling. This effective Fermi surface serves as a mold for the apparition of the half-moon patterns in the equal-time structure factor. Our generic approach enables to extend the notion of half moons to the perturbation of higher-rank Coulomb fields and pinch-line spin liquids. In particular, multi-fold half moons appear when unconventional gauge charges, such as potential fractons, are stabilized in the ground state. Finally, half-moon phases can be tuned across the equivalent of a Lifshitz transition, when the hypersurface manifold changes topology.
@@ I Introduction
Using tools recently developed for the classification of classical spin liquids [5, 67, 68, 16, 18] together with graph-theory methods based on the connectivity matrix [30, 17, 41, 42], we build a generic theory of interacting-cluster systems for a seemingly infinite variety of lattices and clusters in two and three dimensions (or higher). This interaction $\eta$ creates an energetic competition with the zero-divergence field required by the Coulomb gauge theory, and will be our flat-band engineering parameter. We quantify the stability of the flat-band spin liquid until $\eta$ reaches a critical value, where the energetic competition ultimately brings the lowest dispersive band below the flat bands. This crossing selects a hypersurface of parent-system states that are promoted to the ground state of the interacting system, defining an interaction-tunable effective Fermi level despite the absence of particle filling in spin systems. At this point, pinch points transform into half moons, the signature of a spiral spin liquid where the magnetic texture of gauge charges is stabilized in the ground state. Then we extend the notion of half moons to higher-rank U(1) gauge fields, including models with pinch lines. Since multifold half moons come from the dispersive band of a parent rankn-n Coulomb field, they represent the signature of rankn-n gauge charges, such as fractons, in the ground state. Finally, we show how half-moon phases can be adiabatically tuned across the equivalent of a Lifshitz transition separating distinct spiral spin liquids.
@@ II Cluster Hamiltonians
2D Lattice Kagome Checkerboard Hexagonal Kagome hexagonal Square octagon Ruby Square kagome Decorated square kagome Octagonal kagome
Lattice Scheme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Premedial Lattice Scheme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
3D Lattice Pyrochlore Octahedral Kagome bipyramidal Quadrupahedral Hyperkagome
Lattice Scheme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ II.2 Flat-band ground state
From now on, we shall work within the Luttinger-Tisza approximation (LTA) [37, 38, 29], where the spin length constraint is only enforced on average over the entire system, $\sum_{i}|\mathbf{S}_{i}|=N$ with $N$ the total number of spins, which are known to be suitable for cluster-system classical spin liquids, see Appendix C. A known limitation of the LTA is that it cannot always properly account for thermal order by disorder [65, 11], i.e. when thermal fluctuations entropically lift the classical ground-state degeneracy and favor magnetic order at very low temperatures. It is, however, possible to estimate when order by disorder is expected or not, i.e. when flat bands are expected to lead to a magnetically disordered spin-liquid ground state (see section II.5). And even if the system ultimately orders at very low temperature $T_{c}$, the results of this paper remain valid at temperatures above $T_{c}$. Indeed, even if the LTA is an approximation, this method is particularly well suited to the study of classical spin liquids whose disordered magnetic texture accommodates well the averaged spin length constraint, as confirmed by many comparison between theory and simulations [56, 14, 41, 5, 68, 16]. And since we shall only consider isotropic exchange coupling here, the spin component $\alpha\in\{x,y,z\}$ is actually irrelevant and will be omitted from now on.
@@ II.3 The connectivity matrix
$\begin{split}h^{v\leftarrow c}_{\mu,X}(\mathbf{q})&=\sum_{j}h^{v\leftarrow c}_{(0,\mu),(j,X)}e^{i(-\mathbf{R}_{j}+\mathbf{r}_{\mu}-\mathbf{r}_{X}^{c})\cdot\mathbf{q}}\\ &=\sum_{k}h^{v\leftarrow c}_{(k,\mu),(0,X)}e^{i(\mathbf{R}_{k}+\mathbf{r}_{\mu}-\mathbf{r}_{X}^{c})\cdot\mathbf{q}},\\ &=\left[h^{c\leftarrow v}(\mathbf{q})\right]^{\dagger}_{\mu,X}\end{split}$ (12)
@@ III.2 Band dispersion of uniform interacting-cluster models
This derivation can be straightforwardly generalized to cluster systems composed of inequivalent cluster types and involving multiple kinds of geometrical intersections between clusters, such as the two variants of the square kagome lattice shown in Table 1. Details of this generalization are presented in Appendix D.
@@ IV.1 General theory of half moon patterns on uniform interacting-cluster systems
Since $\lambda_{ld}(\mathbf{q}^{\star})=0$ by definition of the the band-touching point $\mathbf{q}^{\star}$, $\mathbf{q}^{\star}$ remains a band-touching point for all values of $\zeta$, even when the dispersive band $\Lambda_{ld}$ becomes negative. By analytical continuity of Eq. (37), the manifold $\mathcal{Q}$ typically forms a closed contour in 2D or a closed surface in 3D, in other words a hypersurface, around $\mathbf{q}^{\star}$ in reciprocal space. The evolution of the band structure as $\zeta$ increases can be understood visually by noting that the dispersion relation in Eq. (37) is given by a polynomial $\Lambda(\lambda)$ acting on the band structure of the parent system, as illustrated in Fig. 3. The spectrum of the interacting system is therefore entirely determined by the behavior of $\Lambda(\lambda)$ over the spectrum of the parent system. In particular, the minimum $\lambda_{c}$ of the polynomial $\Lambda(\lambda)$ plays the role of an effective Fermi level. Indeed, in spin systems there is no notion of filling, and the low-temperature behavior is therefore fully determined by the ground state. However, the introduction of cluster–cluster interactions allows parent-system states at energy $\lambda_{c}$ to become ground states of the interacting system. This mechanism enables one to probe the parent-system energy states located around $\lambda_{c}$. In this sense, $\lambda_{c}$ acts as an effective Fermi level that can be continuously tuned by varying the strength of the cluster–cluster interactions controlled by $\zeta$.
Figure 3: Effective Fermi-level illustration. The band structure of an interacting-cluster system can be written as a polynomial function $\Lambda(\lambda)$ of the band structure of the parent cluster system (see Eq. (37)). (a) Since the parent spectrum contains only positive energies, a positive value of the linear coefficient $\alpha-\eta\Xi/\Omega$ results in a simple, non-uniform rescaling of the parent band structure. (b-c) When the linear coefficient $\alpha-\eta\Xi/\Omega$ becomes negative, the polynomial ceases to be strictly increasing, and the minima of the polynomial select the intersections with the parent band structure as the new ground-state manifold. (b) Just beyond the transition, the new ground-state manifold $\mathcal{Q}$ consists of closed contours encircling the contact points of the parent system. (c) When the polynomial minimum $\lambda_{c}$ crosses a saddle point of the parent-system band structure, the topology of the ground-state manifold changes, corresponding to a Lifshitz transition. Note that the flat bands of the parent system (depicted in red), which are located at zero energy, remain pinned at zero energy in the interacting-cluster spectrum, since the polynomial $\Lambda(\lambda)$ has no constant term.
@@
In absence of extensively degenerate flat bands, this half-moon phase is not a traditional spin liquid. But since the manifold $\mathcal{Q}$ is sub-extensive and isomorphic to an (d1)(d-1)-sphere within a given Brillouin zone, the half-moon phase corresponds to a so-called spiral spin liquid [53, 6, 3, 49, 72, 19] with unusual, non-local, spin dynamics [70]. At very low temperatures, beyond the Luttinger–Tisza approximation, thermal order-by-disorder (ObD) may entropically lift the degeneracy of the manifold $\mathcal{Q}$, particularly in $d=3$ dimensions. In $d=2$ dimensions, however, the Hohenberg–Mermin–Wagner theorem forbids the emergence of long-range magnetic order, although the breaking of discrete lattice rotational symmetries remains possible [44]. Such symmetry breaking, potentially associated with a Berezinskii–Kosterlitz–Thouless–type transition in the case of $XY$ spins, can only occur at very low temperatures. At intermediate temperatures, thermal fluctuations are instead expected to restore the effective degeneracy of $\mathcal{Q}$, even slightly away from fine tuning. As a consequence, the half-moon features in the equal-time structure factor are expected to remain robust above the temperature regime where ObD takes place.
In any case, for $\zeta>\zeta_{c}$, the energy minimum of the interacting-cluster Hamiltonian ($\Lambda_{ld}$) comes from the lowest excited dispersive band of the parent cluster Hamiltonian ($\lambda_{ld}$). Since the flat bands support an emergent Coulomb gauge theory, the lowest dispersive band supports a gapless form of the gauge-charge excitations of the Coulomb theory [41, 69]. From a real-space perspective, the appearance of half-moon features therefore signals that these gauge charges are no longer merely thermal excitations but become an integral part of the ground-state manifold. For two-fold half moons, these gauge charges take the form of magnetic monopoles, corresponding to flux defects that locally violate the flux-conservation constraint imposed by $\bm{\mathcal{C}}=0$. Note that for spins of dimension $n_{s}$, there are $n_{s}$ equivalent copies of such scalar gauge charges. In simple systems where the constrainer corresponds to the local cluster magnetization, combinations of the $n_{s}$ types of gauge charges can be interpreted as effective magnetic moments localized on individual clusters. For multifold half moons, the situation is more subtle. In this case, the emergent gauge field is no longer vectorial, and the corresponding gauge charges cannot be interpreted as simple magnetic dipoles. Nevertheless, they remain associated with local violations of a tensorial Gauss law, and thus correspond to higher-rank, spatially structured charge defects in real space. These objects represent generalized gauge charges whose correlations define the characteristic magnetic textures underlying multifold half-moon patterns. The type of flat-band engineering we have introduced via the cluster interaction $\eta$ thus stabilizes magnetic textures where gauge charges have become part of the ground state. At low temperature, fluctuations are restricted to perturbations close to the hypersurface manifold. But at intermediate temperatures, when the flat bands start to be populated, there is a temperature window where a dense phase of gauge charges co-exist with the vacuum of the flat bands. Although this regime may appear reminiscent of a conventional Coulomb phase, the cluster–cluster interaction introduces an additional dominant interaction between gauge charges. This interaction sets a characteristic length scale for their spatial organization, directly related to the inverse radius of the ground-state hypersurface. Because spins are shared between neighboring clusters, it is generally impossible to simultaneously maximize the norms of oppositely oriented neighboring constrainers. This geometric frustration leads to nontrivial charge arrangements in real space, where a maximally charged cluster is typically surrounded by oppositely charged clusters carrying smaller individual charges, as previously observed in Ref. [41]. By varying the parameter $\eta$, one can include more or less of the dispersive band below the flat bands, i.e. qualitatively more or less gauge charges in the low-energy regime.
For models where the thermal order-by-disorder selects a long-range magnetic order at very low temperatures, the ObD transition would correspond to the crystallization of these gauge charges [7, 41].
@@ IV.5.1 Gapless systems
For models with a single cluster type (e.g. checkerboard, octochlore), one has $\lambda(\mathbf{q})\equiv|\mathbf{L}(\mathbf{q})|^{2}$. If the parent system dispersive band $\lambda(\mathbf{q})$ possesses a saddle point, then, as the threshold $\lambda_{c}(\zeta)$ rises past that saddle energy $E_{s}$ (see Eqs. (52,67)), the topology of the ground–state manifold $\mathcal{Q}$ changes, as illustrated in Figs. 3(b,c) and 8(b) for the generalized checkerboard lattice. The key point to understand is that the manifold $\mathcal{Q}$ varies continuously with $\zeta$. For small $\zeta$, the manifold $\mathcal{Q}$ encircles the pinch points as expected for half moons; see the blue contours on the weight $W_{2}(\mathbf{q})$ of the dispersive band (right panel of Fig. 8(a)) and the half moons in the structure factor (left panel).
@@ IV.5.2 Gapped systems
$\Leftrightarrow\zeta$ $\geq$ $\zeta_{\Delta}>\zeta_{c}(\Delta)$
@@ V Discussion
Here we have built a generic theory for multifold half moons, encompassing the evolution of pinch lines into extended surfaces as signatures of spiral spin liquids whose magnetic texture hosts higher-rank gauge charges in the ground state. The manifold of minimum energy, forming a closed contour in reciprocal space, behaves as an effective Fermi surface of the parent noninteracting system and undergoes topological changes at Lifshitz transitions, when the dispersive band of the parent cluster Hamiltonian possesses saddle points. In most, if not all, three-dimensional models, thermal order by disorder will take place at very low temperature and long-range order the system. As was the case on pyrochlore [41], this magnetic ordering can be seen as the crystallization of the gauge charges that populate the dispersive band. In the case of higher rank spin liquids, one might thus be able to observe a crystallization of fractons. When the parent cluster Hamiltonian is gapped, half moons disappear but spiral spin liquids and Lifshitz transitions remain.
The present work is primarily aimed at establishing a generic and unifying theoretical framework for interacting cluster systems, rather than proposing specific microscopic Hamiltonians tailored for immediate experimental realization. In particular, the cluster–cluster interactions required for the emergence of half-moon patterns may effectively generate couplings between sites at relatively long distances, which render the models discussed here challenging to realize in conventional magnetic materials. Nevertheless, cluster-based descriptions are a well-established and physically motivated route for stabilizing classical spin liquids, and the mechanisms identified in this work are expected to apply broadly to any system whose low-energy physics admits an effective cluster formulation.
The present framework offers an infinite variety of theoretical models supporting a dense phase of gapless gauge charges at intermediate temperature, stabilized by cluster interaction. Their non-local dynamics is an open question, and their connection with local momentum vortices found in spiral spin liquids [70] – equivalent to quadrupoles of fractons – is an appealing direction to investigate. The investigation of half moons in quantum spins $S=1/2$ has given interesting physics on kagome [8, 36, 31] and can now be applied to higher-rank and pinch-line spin liquids.
@@ Appendix B Self Consistent Gaussian Approximation
$\begin{split}\mathcal{H}&=\frac{1}{2}\sum_{i,j}H_{i,j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}\\ &=\frac{1}{2N}\sum_{\mathbf{q}}\sum_{\mu,\nu}H_{\mu\nu}(\mathbf{q})\mathbf{S}_{\mu}(\mathbf{q})\cdot\mathbf{S}_{\nu}(-\mathbf{q})\end{split}$ (87)
@@
Appendix C Relevance of the GLT method for clusters systems
For cluster systems, being able the express the Hamiltonian as a sum of squares
$\begin{split}\mathcal{H}&=\sum_{i,j}H_{ij}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}\\ &=\sum_{n,X}|\bm{\mathcal{C}}_{n,X}|^{2}-N_{u.c}\sum_{X}\sum_{i\in X}\left(\gamma_{i}^{X}\right)^{2}\end{split}$ (94)
corresponds in fact to identify that the Langrange multipliers allowing to find a real space physical ground state are $\gamma_{i}^{X}$, as this allows to express the Hamiltonian as a positive definite term which is minimized by configurations satisfying the ground state constraint $\bm{\mathcal{C}}_{n,X}=0\;\forall\;n,X$. Fourier transforming this Hamiltonian leads to
$\begin{split}\mathcal{H}=\sum_{\mathbf{q}}\sum_{X}\sum_{\mu,\nu}\left[h_{\mu}^{X}(\mathbf{q})\mathbf{S}_{\mu}(\mathbf{q})\right]\cdot\left[h_{\nu}^{X}(-\mathbf{q})\mathbf{S}_{\nu}(-\mathbf{q})\right]\\ -\lambda N_{u.c}\sum_{\mu}\alpha_{\mu}^{2}\end{split}$ (95)
where we defined the parameters $\lambda,\alpha_{\mu}$ such that
$\lambda\sum_{\mu}\alpha_{\mu}^{2}=\sum_{X}\sum_{i\in X}\left(\gamma_{i}^{X}\right)^{2}$ (96)
where $h^{X}_{\mu}(\mathbf{q})$ is the $\mu$ component of Fourier transform of the constrainer $\bm{\mathcal{C}}_{0,X}$ as explained in the main text. Under this form the Hamiltonian corresponds to the Generalized Luttinger Tisza (GLT) method [29]. This procedure encodes a soft spin length constraint
$\sum_{n,\mu}\alpha_{\mu}^{2}|\mathbf{S}_{n,\mu}|^{2}=\sum_{\mathbf{q}}\sum_{\mu}\alpha_{\mu}^{2}|\mathbf{S}_{\mu}(\mathbf{q})|^{2}=N_{u.c}\sum_{\mu}\alpha_{\mu}^{2},$ (97)
thanks to a unique Lagrange multiplier $\lambda$, associated with a set of parameters $\alpha_{\mu}$, specific to the different sublattices, that can be cleverly chosen by a analyzing the symmetries of the problem. For a generic system, encoding this soft constraint would lead to a Hamiltonian
$\begin{split}\mathcal{H}=\sum_{\mathbf{q}}\sum_{\mu,\nu}\left(H_{\mu\nu}(\mathbf{q})-\alpha_{\mu}\alpha_{\nu}\delta_{\mu\nu}\right)\mathbf{S}_{\mu}(\mathbf{q})\cdot\mathbf{S}_{\nu}(-\mathbf{q})\\ +\lambda N_{u.c}\sum_{\mu}\alpha_{\mu}^{2}.\end{split}$ (98)
Therefore, solving the ground state would amount to diagonalize the matrix
$F_{\mu\nu}=\frac{1}{\alpha_{\mu}\alpha_{\nu}}H_{\mu\nu}(\mathbf{q})$ (99)
as the configuration associated with its lowest eigenvalue will naturally minimize the Hamiltonian (98). Finally, one would need to check the resulting configuration fulfills the strong condition $|\mathbf{S}_{i}|=1$ to ensure the built state is indeed a physical state.
For cluster systems, however, the story is different. Indeed, because the real-space structure of the Hamiltonian imposes a macroscopic number of ground-state configurations, one should not expect a unique configuration associated with the lowest eigenvalue of $F$, but rather a macroscopic number, since the lowest eigenvalues of $F$ form a flat band in reciprocal space. This macroscopic degeneracy originates from local constraints encoded in the cluster structure. Consequently, for these systems, the appropriate choice of parameters $\alpha_{\mu}$ is given by Eq. (96) and corresponds to selecting parameters such that the matrix $F$ admits flat bands at the bottom of its spectrum. In this setting, checking that a single obtained configuration satisfies the strong constraint $|\mathbf{S}_{i}|=1$ should be replaced by demonstrating that there exists a macroscopic family of superpositions of flat-band states that fulfill this constraint. Since flat bands are generically associated with the existence of compact localized states [57, 68], these can typically be superposed to form physical ground-state configurations, even in systems with inequivalent sublattices. This scenario has been confirmed, for example, by Monte Carlo simulations for the centered pyrochlore lattice [47].
The GLT method aims to determine the ground state of spin systems, but it does not address the excited energy bands. When discussing the upper part of the spectrum of the Hamiltonian matrix, one should therefore refer to the large-$\mathcal{N}$ approximation, in which the strong constraint $|\mathbf{S}_{i}|=1$ is relaxed and the spin degrees of freedom are treated as scalar fields. While, for the reasons discussed above in the context of the GLT, this approximation provides excellent results for studies focusing on the properties of the ground-state manifold, there is no a priori reason to consider it reliable for higher, dispersive energy bands. However, in the particular case of spiral spin liquids discussed in the present manuscript, the ground-state manifold obtained within the large-$\mathcal{N}$ approximation also displays a large degeneracy, suggesting again the existence of superpositions of these states that realize physical ground states. This explains the excellent agreement between Monte Carlo simulations and large-$\mathcal{N}$ predictions in previous observations of half-moon patterns for interacting cluster versions of the kagome and pyrochlore lattices [41]. Since the ground-state manifold realizes a cut of the non-interacting band-system band structure, such an agreement indicates that the band structure obtained within the large-$\mathcal{N}$ approximation is reliable, at least for systems with equivalent sublattices. A complementary way to understand the success of the large-$\mathcal{N}$ approximation for cluster spin liquid systems is that the extensive entropy associated with the massively degenerate ground-state manifold allows for strong thermal fluctuations even at low temperature, thereby washing out small discrepancies between the true ground state and the states predicted within the approximation when considering observables such as the structure factor.
Appendix D General case of non equivalent interacting clusters
The main text derivation of Sec. III.2 can in fact be generalized to the case of cluster systems composed of inequivalent types of clusters and mixing different type of geometrical intersections between clusters, as for example the two versions of the square kagome lattice depicted in Table. 1. This requires to generalize the Hamiltonian (24) to allow for more complex cluster-cluster connectivity, setting it as
$\mathcal{H}=\alpha\sum_{n}|\bm{\mathcal{C}}_{n}|^{2}+2\eta\sum_{\langle m,n\rangle}\Gamma_{mn}\bm{\mathcal{C}}_{m}\cdot\bm{\mathcal{C}}_{n}.$ (100)
This Hamiltonian can also be well expressed in the framework of connectivity matrices, defining the connectivity matrix $h^{c\leftarrow c}$ as a $N_{c}\times N_{c}$ connectivity matrix linking cluster centers to centers of neighboring clusters. Its coefficients $h^{c\leftarrow c}_{mn}$ are equal to $\Gamma_{mn}$ if the clusters $m$ and $n$ are neighboring clusters, and zero otherwise. Now one can choose the cluster-cluster links weights $\Gamma_{mn}$ such that
$h^{c\leftarrow v}h^{v\leftarrow c}=h^{c\leftarrow c}+\Xi I_{N_{c}}.$ (101)
This corresponds in fact to represent the link between a pair of clusters $(m,n)$ as a sum of the different two move paths going from the cluster center-$m$ to a vertex shared by the clusters $m$ and $n$ and then to the center of the cluster $n$. With this construction the Hamiltonian matrix expresses as
$H=\left(\alpha-\eta\Xi\right)H^{(1)}+\eta\left(H^{(1)}\right)^{2}$ (102)
and the band structure of the interaction cluster model can therefore be expressed as
$\Lambda_{\mu}(\mathbf{q})=\left(\alpha-\eta\Xi\right)\lambda_{\mu}(\mathbf{q})+\eta\lambda_{\mu}(\mathbf{q})^{2}.$ (103)
For systems satisfying the set of conditions (i) to (iii) the present choice cluster-cluster interaction would correspond to encapsulate the coefficient $\Omega$ directly within $\eta$.